1. A constant amount of gas occupies 16.2 L at 313 K. What volume would that gas occupy at 336 K?

2. What’s the molarity of a solution containing 48.5 g of MgCl2 (Molar mass = 95.21 g/mol) in 1.75 L volume?

3. 7.215 moles of a gas occupies 105.1 L at a constant pressure. How many moles of that gas could occupy 175.6 L?

4.What volume of a 0.85 M stock solution would be needed to prepare 4.5 L of a solution at 0.30 M concentration?

5. Given [H3O+] = 3.986x10-5 M, find the pH.

6.Given [H3O+] = 4.59x10-3 M, find the [OH-].
Answer #1

1. Initial volume of gas (~) = 16.2L Initial temperature (11) = 313K Final temperature (T2) = 33.6k final volume (^2) = ? froNow, Poh = 11. 66 or log Lov-1 = 18166 log [ou] = - 11.66 ori [OH-] = 10-11.66 = 2.19x 10-12 So, [OH-] = 2.19 8 10 12 M 2.19x